# ]Present address: Sandia National Laboratories, 1515 Eubank Blvd SE, Albuquerque, NM 87123, USA Multi-loop and Multi-axis Atomtronic Sagnac Interferometry

### v1

@@ Abstract
The interference of matter-waves is fundamental to quantum mechanics. The development of laser cooling, magnetic and optical trapping, and evaporation techniques has powered the field of atom interferometry to rapid growth. Atom interferometry is typically done by coherently splitting a cloud of cold (ultracold) atoms into separate paths using light pulses, propagating them along different trajectories, and finally recombining them to generate interference fringes [1, 2, 3, 4, 5]. The phase of the fringe pattern is a direct measure of the phase difference accumulated by the wavepackets along their respective paths. The interferometer space-time geometry can be configured to perform precise measurements of inertial forces, fundamental constants, dark energy searches, gravitational wave detection, general relativity tests, and more [6, 7, 8, 9, 10, 11].
@@
Atom interferometry is implemented in two main settings: with freely falling cold atoms [5, 12, 13, 14], or with ultracold atoms that are fully trapped [15, 16, 17, 18, 19] or guided [20, 21, 22, 23, 24, 25, 19] in smooth optical and/or magnetic potentials. The latter, also known as the atomtronics approach, aims at developing tools to manipulate atomic matter-waves in controlled environments such as waveguides [11]. Our atomtronic Sagnac interferometer operates by splitting, reflecting, and recombining a Bose-Einstein condensate (BEC) using Bragg diffraction from a pulsed standing wave, with the BEC being guided within a linear optical waveguide throughout the process [26]. To enclose an area, the guide is moved outward and back between the first and the last beam splitter pulse. A key distinction from other guided schemes is that we work with a delta-kick collimated [27, 28] large-size BEC containing only a few thousand atoms resulting in significantly decreased mean-field interactions. This enables us to push the interferometer interrogation time to hundreds of milliseconds while maintaining high fringe contrast. A practical approach to enclosing a larger geometrical area is working with multiple smaller loops. Area-enclosing multi- loop guided atom interferometers have been proposed [21, 29] and experimentally studied with freely falling atoms [30, 31]. With access to a longer interrogation time thanks to the delta-kick cooling, small BECs, and fringe phase correction using a classical accelerometer, we are able to increase the number of loops beyond one and extend the Sagnac area by more than one order compared to our previous generation interferometer [26]. In a two-dimensional magnetic trap (time-orbiting potential trap), Beydler et al. have recently achieved a two orbit BEC interferometer enclosing 8 mm2 Sagnac area [32]. A unique and powerful feature of our scheme is that by moving the guide in different planes (keeping the waveguide’s propagation direction normal to gravity), it is easily possible to sense rotation rates about multiple axes without the issue of wavepacket trajectories being impacted by the gravity. This is not straightforward to implement in other guided schemes, where gravity compensation would be needed when the loop plane is tilted away from horizontal [32].
@@
The interferometer sequence begins right after the delta-kick pulse prepares a collimated wavepacket. Fig. 1 explains the sequence of our moving guide interferometer [26]. The guide translation is performed by a one-axis acousto-optic deflector (AOD, IntraAction ATD-274HD6), placed one focal length () behind a cm plano-convex lens. The focus of this lens is aligned to the trap inside the science cell. By changing the RF frequency to the AOD, the beam after the lens is translated by m for 1 MHz change in the RF frequency. The beamsplitters and mirrors of the atom interferometer are implemented through Bragg scattering off a near resonant standing wave formed by retroreflection from a dichroic mirror placed in the guide beam [26]. A double-square intensity pulse is used as a symmetric beam splitter [33] and a Gaussian pulse for the mirror operation on BECs [34]. We employ (or 1.2 cm/s for 87Rb with D2 line) momentum transfer, where nm. The beamsplitter pulse is composed of two square pulses of 22.3 s and 21.2 s with a no-light gap of 37.9 s in between. The Gaussian reflection pulse has a FWHM duration of 44.3 s. The Bragg laser is 12.6 GHz blue-detuned from the transition (see SM2 on Bragg setup). The one-loop interferometer uses a Bragg pulse sequence of beamsplitter-mirror-mirror-beamsplitter. To scan the interferometer phase, the Bragg laser frequency is varied by applying a voltage to the piezo of the laser at the recombination stage (more details in SM3). Right after the recombination pulse, a strong delta-kick pulse is applied that focuses all the three momentum states to a tight spot in 20 ms, at which point the states are separated by 260 m. This final lens drastically improves the signal to noise ratio, which is particularly important given the low atom number and large size of the BECs. We use absorption imaging to count the number of atoms in the three peaks with no time of flight. Final analysis is performed on the absorption images after applying a fringe-removal algorithm [35, 36].
@@
We have installed two navigation grade accelerometers (InnaLabs AI-Q-2110) in the experiment to compute the phase correction. The first one is attached to the retro-reflection mirror mount and the second to the experiment breadboard to which all the optic posts are mounted. The phase correction given by the two accelerometers agree well. There is a factor of roughly two times average contrast improvement after applying the phase correction. This method of accelerometer correction works equally well for a static or moving guide interferometer and depends only on . Fig 2 shows ms one-loop static guide fringes, without (a) and with (b) the accelerometer phase correction. The peak-to-peak fitted fringe contrast increased from 0.28 to 0.66 by applying the correction. Here, the guide axial confinement is weaker than the moving guide cases in the results below.
@@
In our configuration, each wave-packet traverses the whole loop an integer number of times, and thus the interferometer is only sensitive to rotation and not acceleration. The interferometer phase shift from the rotation of the apparatus (or Earth) is , where A is the loop area, is the rotation angular velocity, and is the atomic mass. It is of course critical to identify and minimize any systematic phase errors.
@@
Fig. 3 shows our multi-loop and large-area interferometer experiment results. In all the fringes discussed below, the radial trapping of the guide was 140 Hz. In the first three fringes of Fig. 3, a to c, the guide was moved by 2.7 mm in 50 ms vertically downward and then brought back to the starting position in 50 ms. The inclusion of hold intervals where the guide position was held fixed increased the single-loop time to the values given below. All the fringes presented below are phase corrected from the accelerometer. Fig. 3a shows a contrast fringe for one-loop moving guide interferometer with = 124 ms and enclosed area of 2.9 mm2. The guide was static for 24 ms and moved for 100 ms in total. For the two and three-loop fringes, the guide was static for 22 ms and moved for 100 ms in each loop. Fig. 3b shows a two-loop contrast fringe fringe with = 246 ms and 5.8 mm2 enclosed area. The three-loop interferometer experiment was a direct extension of the two-loop one, see Fig. 3c. We set ms and the computed enclosed area is 8.7 mm2. The fitted fringe contrast is Fig. 3d is a five-loop fringe (contrast ) of total interrogation time of 108 ms. The was set to zero for this five-loop configuration. The guide translation here was smaller, 0.17 mm.
@@
We now discuss the factors that can limit the performance of our interferometer. The fringe underlying contrast is affected by factors such as overlap of the wave-packets at the end, non-optimal Bragg pulse intensity, imperfect overlap of the Bragg and guide beams, experiment table floating condition, the Zeeman state of the atoms, the axial curvature of the guide, and the starting position of the wave-packets with respect to the waist of the guide beam. We do measure a significantly lower underlying fringe contrast when the interferometer is operated with the initial location of the wavepacket loacted away from the guide waist and/or with high axial curvature [41, 43]. The axial trap frequency sets the size of the collimated BEC (to be the ground state size) that can be used in the interferometer or the size oscillates after the collimation lens and degrades the interferometer contrast.
@@
We thank Ceren Uzun for analyzing all of the experimental data, for developing computer simulations which guided the experiment, for designing and implementing the delta-kick cooling protocol, and for contributing to data collection and manuscript preparation. This work was funded by DARPA under the A-PhI program. The initial development of the experimental apparatus was supported by the Office of Naval Research and by the Laboratory Directed Research and Development program of Los Alamos National Laboratory under project number 20180045DR.
@@
We now describe the steps to prepare the atoms in a magnetically insensitive state and then loading them into the guide. With the BEC held in a crossed dipole trap of and , we shine an RF field with its frequency sweeping from 950 kHz 450 kHz in 10 ms, in the presence of a 1 G homogeneous magnetic field. This transfers the atoms from state to other two Zeeman sub-states of . The magnetically sensitive states are filtered from the guide by applying a magnetic quadrupole gradient kick for 10 ms. Finally, we linearly ramp up the guide beam () power to 420 mW, in 200 ms and adiabatically transfer up to 3000 Bose condensed atoms by ramping down () and () beam powers to zero in 200 ms. One experiment cycle takes around 12 s. The guide beam power is kept constant during the interferometer sequence. All the three optical dipole beams are intensity stabilized using photodiodes and closed-loop feedback technique.
We characterize the guide alignment with BECs. After loading the BEC in a crossed dipole trap of guide and the vertical beam and then turning off the vertical beam fast we let the BEC go through shape and center of mass oscillation in the guide. We use long-axis center of mass oscillation of the BEC to precisely tune the guide levelling with respect to the gravity. This method enables us to align the position of vertical dipole beam to the waist of the guide and thus load the atoms at the bottom of the potential. The procedure is iterated between high and low guide axial curvature to decouple guide levelling from the misaligned guide waist. A long probe time of up to 10 s allows us to get rid of any center of mass motion in a sub-hertz confinement trap.
@@
Our Bragg beam setup scheme is shown in Fig. 5. It mainly consists of a laser (New Focus Vortex Plus TLB-6813-P) centered at 780.24 nm wavelength, an amplifier (New Focus TA-7613), an acousto-optic modulator (Isomet 1206C-1), a noise eater (Thorlabs NEL03) , and a few optical elements to focus and guide the laser beam. The laser is 12.6 GHz blue-detuned with respect to the 87Rb transition. The laser is free running with center wavelength at 780.3 nm. The output of the laser is fiber-coupled to the amplifier. The amplifier output goes through a noise eater for power stabilization and allows us to change the power level. Noise eater output is directed to an AOM operating at 80 MHz and its first order diffraction is coupled into the fiber. At the fiber exit, the beam is collimated to a waist diameter of 5.42 mm using a Thorlabs TC25APC-780 triplet collimator. The beam is then reflected with a dichroic mirror towards the science cell. At the other end of the cell, the beam is incident normally on a second dichroic mirror, is reflected back, and thus creating the standing-wave for BEC splitting and reflection operation. The 1064 nm guide beam passes through the dichroic mirrors and is overlapped with the 780 nm standing-wave beam.
@@
The phase of our atom interferometer is, , where is the phase of Bragg pulse with momentum ( in this work). We scan the interferometer phase by changing the Bragg laser frequency at the recombination stage (last pulse). This is done by changing the piezo voltage in the Bragg laser. Changing the laser frequency by at the last pulse changes the phase of the pulse by . Typically, we get a fringe by scanning the piezo voltage from -2 to 2 V (detuning scanned from 11.8 to 13.4 GHz, to ). In each experimental run, the piezo voltage is ramped to a desired value between the second mirror pulse and the recombination pulse. To avoid hysteresis issues, we then ramp the piezo back to -2 V right after the interferometer cycle is over and it stays there. To scan the interferometer fringe over , the laser frequency is changed by 0.64 GHz. This calculation is consistent with the separation between the retro-mirror and trap location of about 12 cm.
We did some tests to characterize the short-term (hours timescale) Bragg laser frequency stability and reproducibility when the piezo voltage is scanned. First, by just measuring the Bragg laser frequency with a wavelength meter. Second, we measured the BEC splitting fidelity at different piezo voltages between -2 and 2V. Lastly, we measured the transmitted intensity of the Bragg beam passing through a Rubidium vapor cell, on a photodiode. By changing the piezo voltage, we can tune the absorption (laser frequency tuned around the D2 line) and thus measured the frequency stability and reproducibility. All of those tests confirmed that the laser frequency (detuning) is stable and reproducible for our current interferometer setting.
@@
Fig. 7 shows the trajectory of the two wave-packets (+ and -) during the interferometer. The red curve is with STA and the blue curve shows if the guide was moved with a constant velocity. It is clear that with STA the wave-packets enclose more area. This is due to the fact that the guide moves slowly in the beginning and at the end and thus the wave-packets move farther along the guide in those regions. This parametric plot is with a short hold time in the beginning and at the end of the guide motion. Fig. 7 shows the loop that was implemented in the one-loop interferometer experiment presented in Fig. 3a ( = 124 ms and 2.9 mm2 enclosed area).
@@
We scan the duration between the first and the second reflection pulse in the interferometer to find the maximum overlap of the wave-packets at the end and thus maximize the fringe contrast. In this way, we keep the interferometer symmetric. The correction to the return time of the wave-packet at the recombination pulse is [41], where is the guide axial trapping frequency and is the time (splitting to reflection). The required change in the mirror to mirror time interval is twice that amount, or, 2. Fig. 8 shows the measured deviation of the mirror to mirror interval from the exact (if in a flat potential) for the wave-packets to overlap at the end, versus . The experimental results shown in the graph are for static guide interferometers, for = 20, 120, 160, 200, and 220 ms. We fit the data to the function , where only is the free parameter. The fitted trap frequency is 0.51 Hz.

### v2

@@ Abstract
The interference of matter-waves is fundamental to quantum mechanics. The development of laser cooling, magnetic and optical trapping, and evaporation techniques has powered the field of atom interferometry to rapid growth. Atom interferometry is typically done by coherently splitting a cloud of cold (ultracold) atoms into separate paths using light pulses, propagating them along different trajectories, and finally recombining them to generate interference fringes [1, 2, 3, 4, 5]. The phase of the fringe pattern is a direct measure of the phase difference accumulated by the wavepackets along their respective paths. The interferometer space-time geometry can be configured to perform precise measurements of inertial forces, fundamental constants, dark energy searches, gravitational wave detection, general relativity tests, and more [6, 7, 8, 9, 10, 11, 12, 13, 14].
@@
Atom interferometry is implemented in two main settings: with freely falling cold and ultracold atoms [5, 15, 16, 17, 18, 19], or with ultracold atoms that are fully trapped [20, 21, 22, 23, 24, 25, 26] or guided [27, 28, 29, 30, 31, 32, 24, 33] in smooth optical and/or magnetic potentials. The latter, also known as the atomtronics approach, aims at developing tools to manipulate atomic matter-waves in controlled environments such as waveguides [11]. The advantage of this approach is that trapping the matter waves makes possible long interrogation times, and hence high sensitivity, in a physics package that can be much smaller than a system that must be large enough to accommodate atoms falling freely under gravity. Our atomtronic Sagnac interferometer operates by splitting, reflecting, and recombining a Bose-Einstein condensate (BEC) using Bragg diffraction from a pulsed standing wave, with the BEC being guided within a linear optical waveguide throughout the process [34]. To enclose an area, the guide is moved outward and back between the first and the last beam splitter pulse. A key distinction from other guided schemes is that we work with a delta-kick collimated [35, 36, 37] large-size BEC containing only a few thousand atoms resulting in significantly decreased mean-field interactions. This enables us to push the interferometer interrogation time to hundreds of milliseconds while maintaining high fringe contrast. A practical approach to enclosing a larger geometrical area is working with multiple smaller loops. Area-enclosing multi- loop guided atom interferometers have been proposed [28, 38] and experimentally studied with freely falling atoms [39, 40]. With access to a longer interrogation time thanks to the delta-kick cooling, small BECs, and fringe phase correction using a classical accelerometer, we are able to increase the number of loops beyond one and extend the Sagnac area by more than one order compared to our previous generation interferometer [34]. In a two-dimensional magnetic trap (time-orbiting potential trap), Beydler et al. have recently achieved a two orbit BEC interferometer enclosing 8 mm2 Sagnac area [41]. A unique and powerful feature of our scheme is that by moving the guide in different planes (keeping the waveguide’s propagation direction normal to gravity), it is easily possible to sense rotation rates about multiple axes without the issue of wavepacket trajectories being impacted by the gravity. This is not straightforward to implement in other guided schemes, where gravity compensation would be needed when the loop plane is tilted away from horizontal [41].
@@
The interferometer sequence begins right after the delta-kick pulse prepares a collimated wavepacket. Fig. 1 explains the sequence of our moving guide interferometer [34]. The guide translation is performed by a one-axis acousto-optic deflector (AOD, IntraAction ATD-274HD6), placed one focal length () behind a cm plano-convex lens. The focus of this lens is aligned to the trap inside the science cell. By changing the RF frequency to the AOD, the beam after the lens is translated by m for 1 MHz change in the RF frequency. The beamsplitters and mirrors of the atom interferometer are implemented through Bragg scattering off a near resonant standing wave formed by retroreflection from a dichroic mirror placed in the guide beam [34]. A double-square intensity pulse is used as a symmetric beam splitter [42] and a Gaussian pulse for the mirror operation on BECs [43]. We measure nearly 100% splitting and reflection pulse efficiency. We employ (or 1.2 cm/s for 87Rb with D2 line) momentum transfer, where nm. The beamsplitter pulse is composed of two square pulses of 22.3 s and 21.2 s with a no-light gap of 37.9 s in between. The Gaussian reflection pulse has a FWHM duration of 44.3 s. The Bragg laser is 12.6 GHz blue-detuned from the transition (see SM2 on Bragg setup). The one-loop interferometer uses a Bragg pulse sequence of beamsplitter-mirror-mirror-beamsplitter. To scan the interferometer phase, the Bragg laser frequency is varied by applying a voltage to the piezo of the laser at the recombination stage (more details in SM3). Right after the recombination pulse, a strong delta-kick pulse is applied that focuses all the three momentum states to a tight spot in 20 ms, at which point the states are separated by 260 m. This final lens drastically improves the signal to noise ratio, which is particularly important given the low atom number and large size of the BECs. We use absorption imaging to count the number of atoms in the three peaks with no time of flight. Final analysis is performed on the absorption images after applying a fringe-removal algorithm [44, 45].
@@
We have installed two navigation grade accelerometers (InnaLabs AI-Q-2110) in the experiment to compute the phase correction. The first one is attached to the retro-reflection mirror mount and the second to the experiment breadboard to which all the optic posts are mounted. The phase correction given by the two accelerometers agree well. There is a factor of roughly two times average contrast improvement after applying the phase correction. This method of accelerometer correction works equally well for a static or moving guide interferometer and depends only on . Fig 2 shows ms one-loop static guide fringes, without (a) and with (b) the accelerometer phase correction. The peak-to-peak fitted fringe contrast increased from 0.23 to 0.6 by applying the correction. Here, the guide axial confinement is weaker than the moving guide cases in the results below.
@@
In our configuration, each wave-packet completes an integer number of full orbits around the Saganc loop, with the final beamsplitter readout occurring at the same point as the initial splitting. As a result, the interferometer is sensitive to rotation and but not to constant acceleration. The interferometer phase shift from the rotation of the apparatus (or Earth) is , where A is the loop area, is the rotation angular velocity, and is the atomic mass. It is of course critical to identify and minimize any systematic phase errors.
@@
Fig. 3 shows our multi-loop and large-area interferometer experiment results. In all the fringes discussed below, the radial trapping of the guide was 140 Hz. In the first three fringes of Fig. 3, a to c, the guide was moved by 2.7 mm in 50 ms vertically downward and then brought back to the starting position in 50 ms. The inclusion of hold intervals where the guide position was held fixed increased the single-loop time to the values given below. All the fringes presented below are phase corrected from the accelerometer. Fig. 3a shows a contrast fringe for one-loop moving guide interferometer with = 124 ms and enclosed area of 2.9 mm2. The guide was static for 24 ms and moved for 100 ms in total. For the two and three-loop fringes, the guide was static for 22 ms and moved for 100 ms in each loop. Fig. 3b shows a two-loop contrast fringe with = 246 ms and 5.8 mm2 enclosed area. The three-loop interferometer experiment was a direct extension of the two-loop one, see Fig. 3c. We set ms and the computed enclosed area is 8.7 mm2. The fitted fringe contrast is Fig. 3d is a five-loop fringe (contrast ) of total interrogation time of 108 ms. The was set to zero for this five-loop configuration. The guide translation here was smaller, 0.17 mm.
@@
We now discuss the factors that can limit the performance of our interferometer. The fringe underlying contrast is affected by factors such as overlap of the wave-packets at the end, non-optimal Bragg pulse intensity, imperfect overlap of the Bragg and guide beams, experiment table floating condition, the Zeeman state of the atoms, the axial curvature of the guide, and the starting position of the wave-packets with respect to the waist of the guide beam. We do measure a significantly lower underlying fringe contrast when the interferometer is operated with the initial location of the wavepacket located away from the guide waist and/or with high axial curvature [50, 52]. The axial trap frequency sets the size of the collimated BEC (to be the ground state size) that should ideally be used in the interferometer. Otherwise, the long-axis size after the collimation lens oscillates and degrades the interferometer contrast.
@@
The data that support the findings of this study are available from the corresponding author upon reasonable request.
The authors declare no competing interests.
This work was funded by DARPA under the A-PhI program. The initial development of the experimental apparatus was supported by the Office of Naval Research and by the Laboratory Directed Research and Development program of Los Alamos National Laboratory under project number 20180045DR.
@@
We now describe the steps to prepare the atoms in a magnetically insensitive state and then loading them into the guide. With the BEC held in a crossed dipole trap of and , we shine an RF field with its frequency sweeping from 950 kHz 450 kHz in 10 ms, in the presence of a 1 G homogeneous magnetic field. This transfers the atoms from state to other two Zeeman sub-states of . The magnetically sensitive states are filtered from the guide by applying a magnetic quadrupole gradient kick for 10 ms. Finally, we linearly ramp up the guide beam () power to 420 mW, in 200 ms and adiabatically transfer up to 3000 Bose condensed atoms by ramping down () and () beam powers to zero in 200 ms. One experiment cycle takes around 12 s. The guide beam power is kept constant during the interferometer sequence. All three optical dipole beams are intensity stabilized using photodiodes and closed-loop feedback technique.
We characterize the guide alignment with BECs. After loading the BEC in a crossed dipole trap of guide and the vertical beam, and then turning off the vertical beam fast we let the BEC go through shape and center of mass oscillation in the guide. We use the long-axis center of mass oscillation of the BEC to precisely tune the guide levelling with respect to the gravity. This method enables us to align the position of vertical dipole beam to the waist of the guide and thus load the atoms at the bottom of the potential. The procedure is iterated between high and low guide axial curvature to decouple guide levelling from the misaligned guide waist. A long probe time of up to 10 s allows us to get rid of any center of mass motion in a sub-hertz confinement trap.
@@
Our Bragg beam setup scheme is shown in Fig. 5. It mainly consists of a laser (New Focus Vortex Plus TLB-6813-P) centered at 780.24 nm wavelength, an amplifier (New Focus TA-7613), an acousto-optic modulator (Isomet 1206C-1), a noise eater (Thorlabs NEL03), and a few optical elements to focus and guide the laser beam. The laser is 12.6 GHz blue-detuned with respect to the 87Rb transition. The laser is free-running with a center wavelength at 780.3 nm. The output of the laser is fiber-coupled to the amplifier. The amplifier output goes through a noise eater for power stabilization and allows us to change the power level. Noise eater output is directed to an AOM operating at 80 MHz and its first order diffraction is coupled into the fiber. At the fiber exit, the beam is collimated to a waist diameter of 5.42 mm using a Thorlabs TC25APC-780 triplet collimator. The beam is then reflected with a dichroic mirror towards the science cell. At the other end of the cell, the beam is incident normally on a second dichroic mirror, is reflected, and thus creates the standing-wave for the BEC splitting and reflection operation. The 1064 nm guide beam passes through the dichroic mirrors and is overlapped with the 780 nm standing-wave beam.
@@
The phase of our atom interferometer is, , where is the phase of Bragg pulse with momentum ( in this work). We scan the interferometer phase by changing the Bragg laser frequency at the recombination stage (last pulse). This is done by changing the piezo voltage in the Bragg laser. Changing the laser frequency by at the last pulse changes the phase of the pulse by , where is the distance between the BEC and the retroreflecting mirror. Typically, we get a fringe by scanning the piezo voltage from -2 to 2 V (detuning scanned from 11.8 to 13.4 GHz, to ). In each experimental run, the piezo voltage is ramped to a desired value between the second mirror pulse and the recombination pulse. To avoid hysteresis issues, we then ramp the piezo back to -2 V right after the interferometer cycle is over and it stays there. To scan the interferometer fringe over , the laser frequency is changed by 0.64 GHz. This calculation is consistent with the separation between the retro-mirror and trap location of about 12 cm.
We did some tests to characterize the long-term (hours timescale) Bragg laser frequency stability and reproducibility when the piezo voltage is scanned. First, by just measuring the Bragg laser frequency with a wavelength meter. Second, we measured the BEC splitting fidelity at different piezo voltages between -2 and 2V. Lastly, we measured the transmitted intensity of the Bragg beam passing through a Rubidium vapor cell, on a photodiode. By changing the piezo voltage, we can tune the absorption (laser frequency tuned around the D2 line) and thus measure the frequency stability and reproducibility. All of those tests confirmed that the laser frequency (detuning) is stable at about 100 MHz level and reproducible for our current interferometer setting.
@@
Fig. 7 shows the trajectory of the two wave-packets (+ and -) during the interferometer. The red curve is with STA, and the blue curve shows if the guide was moved with a constant velocity. It is clear that with STA the wave-packets enclose more area. This is due to the fact that the guide moves slowly in the beginning and at the end, and thus the wave-packets move farther along the guide in those regions. This parametric plot is with a short hold time in the beginning and at the end of the guide motion. Fig. 7 shows the loop that was implemented in the one-loop interferometer experiment presented in Fig. 3a ( = 124 ms and 2.9 mm2 enclosed area).
@@
We scan the duration between the first and the second reflection pulse in the interferometer to find the maximum overlap of the wave-packets at the end and thus maximize the fringe contrast. In this way, we keep the interferometer symmetric. The correction to the return time of the wave-packet at the recombination pulse is [50], where is the guide axial trapping frequency and is the time (splitting to reflection). The required change in the mirror to mirror time interval is twice that amount, or 2. Fig. 8 shows the measured deviation of the mirror to mirror interval from the exact (if in a flat potential) for the wave-packets to overlap at the end, versus . The experimental results shown in the graph are for static guide interferometers, for = 20, 120, 160, 200, and 220 ms. We fit the data to the function , where only is the free parameter. The fitted trap frequency is 0.51 Hz.